OpenStudy (sbrummitt):

The first side of a triangle measures 5 in. less than the second side, the third side is 3 in. more than the first side, and the perimeter is 17 in. How long is the third side? If s represents the length of the second side, which of the following represents the length of the third side?

OpenStudy (sbrummitt):

ANYBODY

OpenStudy (jiteshmeghwal9):

If s be the second side => first side= s-5 third side = (s-5)+3 perimeter= first side + second side + third side =17

OpenStudy (jiteshmeghwal9):

now u may find s & the length (in inches) of the third side.

OpenStudy (sbrummitt):

thank you.